# Non-commutativity as a Universal Characterization for Enhanced Quantum Metrology

### v1

@@ Abstract
The parameter of interest is the average displacement $\overline{x}=\sum_{j=1}^{N}x_{j}/N$, with an auxiliary squeezing strength $\overline{\xi}=\sum_{i=1}^{N}\xi_{i}/N$. The total unitary evolution is given by $\hat{U}_{x}=\prod_{j=1}^{N}\hat{D}_{x_{j}}\prod_{i=1}^{N}\hat{S}_{\xi_{i}}$, where $\hat{D}_{x_{j}}=e^{-ix_{j}\hat{P}}$ and $\hat{S}_{\xi_{i}}=e^{-i\frac{\xi_{i}}{2}((\hat{a}^{\dagger})^{2}+\hat{a}^{2})}$. Owing to the non-commutative structure between $\hat{H}_{g}$ and $\hat{H}_{\lambda}$, this evolution exhibits an infinite nilpotency index $\mathcal{K}\to\infty$, as evidenced by the unbounded sequence of nested commutators $(\mathrm{ad}_{\hat{H}_{g}})^{n}(\hat{H}_{\lambda})\neq 0$ for all $n\in\mathbb{N}$. This algebraic structure gives rise to the local generator of the form $\hat{h}_{\overline{x}}=-N\sinh(N\overline{\xi})\hat{X}+N\cosh(N\overline{\xi})\hat{P}$. The corresponding QFI reads
@@
For an initial coherent state $|\alpha\rangle$, the QFI simplifies to $2N^{2}\cosh(2N\overline{\xi})$, demonstrating an exponential scaling $F_{\overline{x}}\simeq N^{2}e^{2N}$ when $\overline{\xi}>0$, as shown in Fig. 3. This exponential enhancement arises from the non-commutative nature of the displacement and squeezing operators. Moreover, the final state $|\Psi_{\overline{x}}\rangle=\hat{U}_{x}|\alpha\rangle$ remains Gaussian, enabling efficient readout via quadrature measurements that nearly saturate the QCRB. The generalized quadrature observable is defined as $\hat{Q}=\hat{X}\cos\theta+\hat{P}\sin\theta$, and its variance in the final state is given by
@@
The encoding processes discussed above can be implemented in a programmable loop-based optical processor capable of executing sequential Gaussian gates. As demonstrated in recent experiments, such platforms support the repeated application of programmable single-mode and multi-step squeezing gates, offering a viable physical system for realizing the proposed non-commutative sensing protocols [47, 48]. Notably, despite deeper non-commutative structures generally requiring higher circuit depth, more demanding control, and exhibiting stronger vulnerability to decoherence, the infinite-$\mathcal{K}$ limit can be naturally attained in specific platforms.
Conclusion.—To summarize, we present a quantum metrology framework harnessing non-commutative operator sequences to surpass standard precision limits. By introducing the nilpotency index $\mathcal{K}$ as a universal characterization to quantify the non-commutativity in the encoding dynamics, we derive scaling laws governing quantum advantage. For finite $\mathcal{K}$, the RMSE generically scales as $N^{-(1+\mathcal{K})}$—exceeding the conventional Heisenberg scaling without requiring ICO—provided the $\mathcal{K}$-th order nested commutator remains non-constant. When this commutator degenerates to a constant, ICO becomes essential to restore optimal scaling. When $\mathcal{K}\to\infty$, we show that an exponential enhancement $\Delta\lambda\sim N^{-1}e^{-N}$ can be achieved for a broad class of initial states.
Moving forward, the quantum-enhanced metrology in our protocol arises from the Lie algebraic relations among the operators. While the present analysis focuses on single-parameter estimation, the proposed scheme can be naturally extended to multi-parameter scenarios. In such settings, the RMSE for composite observables, such as $A:=\overline{\lambda}\overline{g}$, is fundamentally related to the two achievable scaling exponents, especially the worse one [44]. Beyond the sequential encoding protocols, the Lie algebraic structure of non-commutative operators can also be leveraged in critical metrology to surpass quantum speed limits at the criticality [40]. This opens an interesting direction for exploring the interplay between non-commutativity and criticality in enhancing sensing performance.
Acknowledgement.—This work is supported by the National Natural Science Foundation of China (Grants No. 12125402, No. 12350006, No. 62575232), Beijing Natural Science Foundation (Grant No. Z240007), and the Innovation Program for Quantum Science and Technology (Grants No. 2024ZD0302401, No. 2021ZD0301500). Y.X. acknowledges support by the National Cryptologic Science Fund of China (Grant No. 2025NCSF02048).

### v2

@@ Abstract
The parameter of interest is the average displacement $\overline{x}=\sum_{j=1}^{N}x_{j}/N$, with an auxiliary squeezing strength $\overline{\xi}=\sum_{i=1}^{N}\xi_{i}/N$. The total unitary evolution is given by $\hat{U}_{x}=\prod_{j=1}^{N}\hat{D}_{x_{j}}\prod_{i=1}^{N}\hat{S}_{\xi_{i}}$, where $\hat{D}_{x_{j}}=e^{-ix_{j}\hat{P}}$ and $\hat{S}_{\xi_{i}}=e^{-i\frac{\xi_{i}}{2}((\hat{a}^{\dagger})^{2}+\hat{a}^{2})}$. Owing to the noncommutative structure between $\hat{H}_{g}$ and $\hat{H}_{\lambda}$, this evolution exhibits an infinite nilpotency index $\mathcal{K}\to\infty$, as evidenced by the unbounded sequence of nested commutators $(\mathrm{ad}_{\hat{H}_{g}})^{n}(\hat{H}_{\lambda})\neq 0$ for all $n\in\mathbb{N}$. This algebraic structure gives rise to the local generator of the form $\hat{h}_{\overline{x}}=-N\sinh(N\overline{\xi})\hat{X}+N\cosh(N\overline{\xi})\hat{P}$. The corresponding QFI reads
@@
For an initial coherent state $|\alpha\rangle$, the QFI simplifies to $2N^{2}\cosh(2N\overline{\xi})$, demonstrating an exponential scaling $F_{\overline{x}}\simeq N^{2}e^{2N}$ when $\overline{\xi}>0$, as shown in Fig. 3. This exponential enhancement arises from the noncommutative nature of the displacement and squeezing operators. Moreover, the final state $|\Psi_{\overline{x}}\rangle=\hat{U}_{x}|\alpha\rangle$ remains Gaussian, enabling efficient readout via quadrature measurements that nearly saturate the QCRB. The generalized quadrature observable is defined as $\hat{Q}=\hat{X}\cos\theta+\hat{P}\sin\theta$, and its variance in the final state is given by
@@
The encoding processes discussed above can be implemented in a programmable loop-based optical processor capable of executing sequential Gaussian gates. As demonstrated in recent experiments, such platforms support the repeated application of programmable single-mode and multi-step squeezing gates, offering a viable physical system for realizing the proposed noncommutative sensing protocols [Takeda2021, Takeda2025]. Notably, despite deeper noncommutative structures generally requiring higher circuit depth, more demanding control, and exhibiting stronger vulnerability to decoherence, the infinite-$\mathcal{K}$ limit can be naturally attained in specific platforms.
Conclusion.—To summarize, we present a quantum metrology framework harnessing noncommutative operator sequences to surpass standard precision limits. By introducing the nilpotency index $\mathcal{K}$ as a universal characterization to quantify the noncommutativity in the encoding dynamics, we derive scaling laws governing quantum advantage. For finite $\mathcal{K}$, the RMSE generically scales as $N^{-(1+\mathcal{K})}$—exceeding the conventional Heisenberg scaling without requiring ICO—provided the $\mathcal{K}$-th order nested commutator remains nonconstant. When this commutator degenerates to a constant, ICO becomes essential to restore optimal scaling. When $\mathcal{K}\to\infty$, we show that an exponential enhancement $\Delta\lambda\sim N^{-1}e^{-N}$ can be achieved for a broad class of initial states.
Moving forward, the quantum-enhanced metrology in our protocol arises from the Lie algebraic relations among the operators. While the present analysis focuses on single-parameter estimation, the proposed scheme can be naturally extended to multi-parameter scenarios. In such settings, the RMSE for composite observables, such as $A:=\overline{\lambda}\overline{g}$, is fundamentally related to the two achievable scaling exponents, especially the worse one [sm]. Beyond the sequential encoding protocols, the Lie algebraic structure of noncommutative operators can also be leveraged in critical metrology to surpass quantum speed limits at the criticality [Cai2021]. This opens an interesting direction for exploring the interplay between noncommutativity and criticality in enhancing sensing performance.
Acknowledgement.—This work was supported by Beijing Natural Science Foundation (Grant No. Z240007), National Natural Science Foundation of China (Grants No. 12125402, No. 12534016, No. 12350006, No. 62575232), and Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2024ZD0302401 and No. 2021ZD0301500). Y.X. acknowledges support by the National Cryptologic Science Fund of China (Grant No. 2025NCSF02048).
Data availability.—The data that support the findings of this article are openly available [data].